# Countering threats to national security posed by AI systems through an incident regime

### v2

@@ Abstract
Recent progress in AI capabilities has heightened concerns that AI systems could pose a threat to national security, for example, by making it easier for malicious actors to perform cyberattacks on critical national infrastructure, or through loss of control of autonomous AI systems. In parallel, federal legislators in the US have proposed nascent ’AI incident regimes’ to identify and counter similar threats. In this paper, we consolidate these two trends and present a timely proposal for a legally mandated post-deployment AI incident regime that aims to counter potential national security threats from AI systems. We start the paper by introducing the concept of ’security-critical’ to describe doctors that pose extreme risks to national security, before arguing that ’security-critical’ describes civilian nuclear power, aviation, life science dual-use research of concern, and frontier AI development. We then present in detail our AI incident regime proposal, justifying each component of the proposal by demonstrating its similarity to US domestic incident regimes in other ’security-critical’ sectors. Finally, we sketch a hypothetical scenario where our proposed AI incident regime deals with an AI cyber incident. Our proposed AI incident regime is split into three phases. The first phase revolves around a novel operationalization of what counts as an ’AI incident’ and we suggest that AI providers must create a ’national security case’ before deploying a frontier AI system. The second and third phases spell out that AI providers should notify a government agency about incidents, and that the government agency should be involved in amending AI providers’ security and safety procedures, in order to counter future threats to national security.
@@ 1 Introduction
While current AI systems likely do not pose threats to national security (Bengio et al., 2025), recently there has been fast progress in the dangerous capabilities that AI systems possess: OpenAI denoted its frontier system released in July 2024 (GPT-4o) as posing a ‘low’ cyber and CBRN risk (Chemical, Biological, Radiological and Nuclear) (OpenAI, 2024a), but its frontier system only 7 months later (o1, released in December 2024) was already designated ‘medium’ risk on both these categories (OpenAI, 2024b). It is not clear if future AI systems will maintain this trend, but in light of recent progress it may be prudent for nation states to build up capacity both to track the national security threats that AI systems pose, and to execute countermeasures to neutralise these threats. Against this backdrop of growing AI capabilities, US federal legislators have started proposing nascent variations of ‘AI incident reporting regimes’ (118th Congress, 2024; Ross, 2024).
@@
We combine these two strands – increased government interest in both the national security risks posed by AI systems, and in AI incident regimes – to focus on the role that a legally mandated post-public deployment AI incident regime could play in giving a government agency oversight into national security threats posed by AI systems. Our ambition is to enable a government agency to maintain comprehensive awareness of AI threats and rapidly counter any resulting risks to national security.
@@
In this paper, we put forward a proposal for an AI incident regime that will help to counter threats to national security posed by AI systems. We start the paper by introducing a novel term – ‘security-critical’ – to describe sectors that pose extreme risks to national security, before arguing that civilian nuclear power, aviation, and life sciences dual-use research of concern, and frontier AI development, should all be considered ‘security-critical’ sectors..We then spend much of the paper detailing our proposal for an AI incident regime that aims to counter national security threats from AI systems. We justify each component of our AI incident proposal by demonstrating that it mirrors existing regimes in security-critical industries, specifically US domestic incident regimes in civilian aviation, life sciences dual-use research of concern (e.g. genetic engineering) and nuclear power.
@@
Informed by this comparison to other high-risk areas, our proposed AI incident regime consists of three phases. First, a ‘Preparatory Phase, which details steps that AI providers should take in order to define which events count as ‘incidents’ ahead of time; then the ‘Rapid Response Phase’, in which AI providers promptly notify a government agency about incidents, and this government agency taking steps to contain the incidents; and finally the ‘Hardening Defenses Phase’, wherein AI providers improve their security and safety procedures in order to counter future threats to national security. We end our paper with a sketch of how our proposal would deal with a hypothetical spear-phishing attack.
@@
In summary, the core contribution of this paper is a proposal for a national security-focused AI incident regime [^1] We note that all references to an AI incident regime in this paper are tied to an AI incident regime specifically for national security threats, rather than other harms which has three phases, with each component touching on either AI providers or the government agency :
@@
      ‘Incident’ should be operationalised as ‘any event which weakens a claim made in a national security case for the AI system’
@@
      Providers of frontier AI systems must notify a government agency immediately, and no later than 24 hours, after they discover that an incident has occurred  [^2] We define frontier AI sytems as ’highly capable general-purpose AI models or systems that can perform a wide variety of tasks and match or exceed the capabilities present in the most advanced models’ (UK Government, 2024). We believe targeting frontier models ensures the most severe threats to national security can be rapidly identified and neutralised, while still enabling broad economic benefits from widespread adoption of less advanced AI systems.
@@
      If appropriate, the government agency shall coordinate an emergency containment to limit harm caused by the incident
@@
      The government agency has the ‘intelligence-gathering’ authority to access information that allows them to investigate the cause of an incidents (consisting for example, of extensive blackbox access, some white-box access, and relevant documentation and technical results [^3] Note that it is not currently possible to use interpretability tools to reliably obtain a ‘decision rationale’ that ‘explains’ an AI systems output. Instead, we are recommending that a government agency has some white-box access to the AI system they are investigating, in order to speed up their blackbox investigation).
@@
Our proposal for an AI incident regime is limited in scope: we detail an AI incident regime focussed purely on countering the most extreme threats to national security (see Section 2 for more justification) given the threat to people’s are to people’s lives and livelihoods (Gaibulloev & Sandler, 2019; Iftikhar, 2024). Further, our proposal is designed to be agile, imposing minimal burden on providers of AI systems that do not pose threats to national security – however, if AI systems start to possess dangerous capabilities, the regime ratchets upwards, offering ample intelligence-gathering and security-strengthening capabilities to a government agency and thus helping to counter potential national security threats from AI systems.
@@
2 Nuclear power, aviation, life sciences, and frontier AI development are all ‘security-critical’ sectors
@@
In this section, we coin the term ’security-critical’, to describe sectors that pose extreme risks to national security. We argue that AI, and all three of our selected case study sectors – nuclear power, aviation, and life sciences DURC – are ‘security-critical’ sectors. This leads us to the claim that incident regimes in these security-critical industries are likely good templates for the design of an AI incident regime: a claim we will use later on to justify each component of our AI incident regime proposal. However, we spend most of this section demonstrating that there is a similarity in the threats to national security posed by AI systems and threats posed in nuclear power, aviation, and life sciences DURC, in order to argue that it is natural to bucket all of these sectors together as ‘security-critical’. We make this argument by first detailing the potential extreme threats posed in nuclear power, aviation, and life sciences DURC, and then detailing the potential extreme threats posed by AI systems, showing that in the worst-case scenarios there could be (at least) thousands of fatalities in a short space of time – making them all ‘security-critical’.
@@
We start by detailing the most extreme threats posed in nuclear power, aviation, and life sciences, demonstrating that these can lead to (at least) thousands of fatalities within months. For example, a 1975 NRC report on extreme risks posed by nuclear power, offers an upper bound of the worst scenario, at 3000 fatalities ‘in a short time period after the incident’ and up to 60,000 indirect fatalities from cancer (Rasmussen, 1975). Turning to historical precedents in nuclear power, the Chernobyl incident led to the deaths of 2 workers within hours of the plant meltdown and 22 within the following month due to the exposure to radiation (United Nations Scientific Committee, 2008), while one estimate from the International Atomic Energy Authority put the indirect number of fatalities due to cancer at 4000 (The Chernobyl Forum, 2006). Looking at historical incidents in aviation, the 1977 Tenerife collision between two aircraft resulted in the deaths of 583 people (Register-Guard, 1977), while the 9/11 attacks resulting in around 3,000 casualties (Encyclopedia Britannica, 2024). Meanwhile, in life sciences, one paper estimates that a potential pandemic pathogen escaping from a lab could lead to tens of millions of fatalities (Klotz & Sylvester, 2014). Historical precedent is demonstrated by events such as the 1979 anthrax lab leak in the Soviet Union led to 68 fatalities within 6 weeks of the leak occurring (Meselson et al., 1994). Meanwhile, Covid-19 is estimated to have resulted in over 7 million casualties as of January 2025 (World Health Organization, 2025), giving some sense of the possible worst-case consequences of a life sciences DURC lab leak.
@@
Now, we demonstrate that the most extreme national security threats from AI could also lead to at least thousands of fatalities within months of systems being deployed, by detailing some of the extreme threats to national security that experts have cautioned about. We then explain that AI systems do not currently pose such threats, but note that if recent trends in AI capability growth continue, such threats will arise in the near future.
@@
First, we detail the worst-case national security threats that AI systems could pose. For example, some experts worry that AI systems could uplift the ability of malicious actors to create bioweapons, (Drexel & Withers, 2024; Reuters, 2023), which could cause a pandemic and lead to fatalities within weeks of initial infection (Baud et al., 2020). Alternatively, malicious actors could use AI systems to help with vulnerability discovery and exploitation for a large-scale cyber attack on critical national infrastructure (Lee et al., 2016; OECD, 2024). Such attacks could, for example, bring down the electricity grid within hours of being executed (Crowdstrike, 2024; Allianz Commercial, 2016). More speculatively, there is also the threat of loss of control of autonomous general-purpose AI systems (Bengio et al., 2025; Hendrycks et al., 2023; OECD, 2024) , in which highly capable AI systems end up misaligned with the intentions of the AI provider (Bengio et al., 2025; Ngo et al., 2025). In this scenario, a misaligned, highly capable AI system could pose a threat to national security by e.g. creating bioweapons or executing a large-scale cyber attack (Bengio et al., 2024; Hendrycks et al., 2023).
@@
Next, we note a worrying trend in AI capabilities development which points towards AI systems posing such threats to national security through cyber, bio, or loss of control threats within the next few years. Despite the fact that as of March 2025, publicly deployed AI systems do not appear to pose much danger (Bengio et al., 2025; Christopher A. Mouton, 2023), recent growth in general AI capabilities has been fast (Maslej et al., 2024; Roser, 2022), and more recently there has been growth in capabilities that pose threats to national security. Regarding the former, Anthropic CEO Dario Amodei has said that within 3 years we will have a ‘country of geniuses in a datacenter’ (Amodei, 2024) and that ‘AI could surpass almost all humans at almost everything’ (Edwards, 2025). Further, a recent report co-authored by 96 world-leading AI experts includes a Chair’s note from Yoshua Bengio claiming that recent evidence points towards ‘the pace of advances in AI capabilities … remain[ing] high or even accelerat[ing]’ (Bengio et al., 2025). Against this backdrop of general AI capabilities increasing, there has also been fast progress in dangerous capabilities that pose threats to national security: a leading AI system from August 2024 (GPT-4o) was rated as posing a ”Low” risk on all of these domains(OpenAI, 2024a) while just 7 months later a frontier system was rated as posing a Medium risk on all these domains (OpenAI’s Deep Research, released in January 2025) (OpenAI, 2025). If this trend is maintained, AI systems will soon pose threats to national security on-par with those from nuclear power, aviation and life sciences DURC.
@@
Hence in this sub-section we demonstrated that AI systems may soon pose national security threats similar to those in nuclear power, life sciences DURC, and aviation, involving thousands of fatalities in a short space of time. Given the similarity in threats, we claim that incident regimes in these security-critical sectors are promising templates for an AI incident regime – and this is a claim that will underpin each part of our AI incident regime proposal.
3 AI Incident Regime Proposal
@@
Our novel proposal for an AI incident regime consists of three phases, each backed up by our review of incident regimes in security-critical sectors: a ‘Preparatory Phase’ where AI providers pre-determine ahead of time what events count as incidents; a ‘Rapid Response Phase’ wherein AI providers promptly notify a government agency about incidents, and the government agency then contains the incident; and a ‘Hardening Defenses Phase’, in which the government agency takes steps to counter future national security threats.
@@
We take most of this Section (3.1-3.3) to articulate in detail each of these phases of our AI incident regime proposal. We justify in turn each component of our proposed AI incident regime by demonstrating that the component mirrors incident regimes in other security-critical sectors. In particular, we make the comparison between our AI incident regime proposal, and: in nuclear power, the incident regime run by the Nuclear Regulatory Commission (NRC); in life sciences the National Institute for Health’s (NIH) incident regime on genetic engineering and the Federal Select Agents Program (FSAP) incident regime on dangerous biological agents; and in aviation, the National Transportation Safety Board’s (NTSB) incident investigations and the Federal Aviation Authority’s (FAA) post-incident rule-making. We then end the paper with a hypothetical scenario tying our blueprint together in Section 4. Before detailing each phase of our AI incident regime, we present a short summary, detailing for each phase the goal, and how we expect AI providers or the government agency to act [^4] The components in our incident regime proposal should not be considered sufficient for an AI incident regime to adequately counter potential threats to national security. For example, in order for the incident regime to be successful, it may be important to set up infrastructure to allow consumers or downstream developers to report incidents (in addition to AI providers), in order to ensure that the regime comprehensively catches incidents. :
@@
      AI providers should report as an incident any event which weakens a claim made in a ‘national security case’ for the AI system
@@
      Providers of frontier AI systems must notify a government agency immediately, and no later than 24 hours, after they discover that an incident has occurred  [^5] We define frontier AI sytems as ’highly capable general-purpose AI models or systems that can perform a wide variety of tasks and match or exceed the capabilities present in the most advanced models’ (UK Government, 2024). We believe targeting frontier models ensures the most severe threats to national security can be rapidly identified and neutralised, while still enabling broad economic benefits from widespread adoption of less advanced AI systems.
@@
      If appropriate, the government agency will coordinate an emergency containment to limit harm caused by the incident
@@
      The government agency has the ‘intelligence-gathering’ authority to perform on incidents through extensive access to the potentially compromised AI system (consisting for example, of extensive blackbox access, some white-box access, and relevant documentation and technical results [^6] Note that it is not currently possible to use interpretability tools to reliably obtain a ‘decision rationale’ that ‘explains’ an AI systems output. Instead, we are recommending that a government agency has some white-box access to the AI system they are investigating, in order to speed up their blackbox investigation).
@@
      A government agency has ‘security-strengthening authority’ to make recommendations as to how AI providers’ safety and security infrastructure should be modified, based on the results of their root cause analyses. Under this authority, all relevant AI providers must implement these recommendations, in order to counter future threats to national security.
@@
3.1 ‘Preparatory Phase’: AI providers should create national security safety cases pre-deployment, and should report as an incident ‘any event which weakens a claim made in the national security case’
@@
The first phase of our proposal for an AI incident regime – featuring a novel operationalisation of ‘AI incident’ – consists of two preparatory steps that AI providers must take in order to quickly report an incident when it arises: (i) AI providers must create national security cases for their frontier AI systems before deploying them publicly and (ii) AI providers must report as an incident ‘any event which weakens the evidence for a claim or argument chain made in an AI system’s national security case.’ We start this section by detailing how an AI provider could comply with each of these with each of these steps, and then go on to justify our proposal for a ‘Preparatory Phase’ with a three-part argument: first, we show that incident regimes in other security-critical sectors have both clear criteria for what counts as an ‘incident’, and include near-misses or precursory harms as incidents. Then we detail how an existing operationalisation of AI incident in the literature fails to fulfill these criteria, before finally showing that our proposal addresses these shortcomings.”
@@
We start by presenting the first obligation for AI providers under the ‘Preparatory Phase’: that AI providers should produce a ‘national security safety case’ before publicly deploying any frontier models [^7] This paper focusses on AI incidents caused by AI systems that have been publicly deployed. In principle, national security cases could be written before public deployment in order to combat to track national security incidents posed by AI providers’ internal deployment of frontier AI systems.. A ‘national security case’ is a security and safety framework adapted from previous frameworks in the literature. It is a document laying out a structured argument making the case that a given AI system does not pose national security risks in public deployment (Ministry of Defence, 2017)– perhaps because it does not pose an unacceptable increase in biorisk, or risk of a large-scale cyber attack, or risk of loss of control of an autonomous AI system [^8] The ‘national security case’ is an adaptation of a previous AI governance framework – AI safety cases – for a national security setting (Balesni et al., 2024; Clymer et al., 2024; Buhl et al., 2024; Goemans et al., 2024; METR, 2024; UK Government, 2024). This structured argument comprises a series of claims, each of which needs to be supported by verifiable evidence about the AI system in question. [^9] It is outside the scope of this paper to make recommendations on the which threats an AI national security case should focus on, and therefore this is a promising question for further research. In addition to the cyber, bio, and loss of control threats already mentioned, some other plausible threats include mass persuasion and threats to the financial system from automated trading). For example, an AI provider might make the high-level claim in their ‘national security case’ that a given AI system does not pose an unacceptable cyber threat. The AI provider might justify this high-level claim as in Figure 1, by a series of more concrete claims that the AI systems cannot help perform certain actors perform certain kinds of cyber attack – for example, claiming that the AI system cannot assist a technical user who is not a cybersecurity expert in discovering and exploiting cyber vulnerabilities in a realistic setting (C3.1). The AI provider then justifies these sub-claims with a body of empirical evidence about the AI system, usually dangerous capability or propensity, evaluation test results – for example, demonstrating through a ‘human uplift’ study that, when paired with an AI systems, a human novice does not improve at cyber vulnerability discovery, compared to doing it by themselves themselves. For more concrete detail on what an AI ‘national security case’ might look like in the cyber domain, see (Goemans et al., 2024).
@@
Figure 1: A visualisation of part of a ‘national security case’ which argues that a given AI system does not pose a cyber threat (taken from (Goemans et al., 2024)).

### v3

# AIs’ threat to national security can be countered through an incident regime

@@ Abstract
Recent progress in AI capabilities has heightened concerns that AI systems could pose a threat to national security, for example, by making it easier for malicious actors to perform cyberattacks on critical national infrastructure or through loss of control of autonomous AI systems. In parallel, federal legislators in the U.S. have proposed nascent ‘AI incident regimes’ to identify and counter similar threats. In this paper, we consolidate these two trends and present a timely proposal for a legally mandated post-deployment AI incident regime that aims to counter potential national security threats from AI systems. We start the paper by introducing the concept of ‘security-critical’ to describe doctors that pose extreme risks to national security before arguing that ’security-critical’ describes civilian nuclear power, aviation, life science dual-use research of concern, and frontier AI development. We then present in detail our AI incident regime proposal, justifying each component of the proposal by demonstrating its similarity to U.S. domestic incident regimes in other ‘security-critical’ sectors. Finally, we sketch a hypothetical scenario where our proposed AI incident regime deals with an AI cyber incident. Our proposed AI incident regime is split into three phases. The first phase revolves around a novel operationalization of what counts as an ‘AI incident,’ and we suggest that AI providers must create a ‘national security case’ before deploying a frontier AI system. The second and third phases spell out that AI providers should notify a government agency about incidents and that the government agency should be involved in amending AI providers’ security and safety procedures in order to counter future threats to national security.
@@ 1 Introduction
While current AI systems likely do not pose threats to national security (Bengio et al., 2025), recently there has been fast progress in the dangerous capabilities that AI systems possess: OpenAI denoted its frontier system released in July 2024 (GPT-4o) as posing a ‘low’ cyber and CBRN risk (Chemical, Biological, Radiological and Nuclear) (OpenAI, 2024a), but its frontier system only 7 months later (o1, released in December 2024) was already designated ‘medium’ risk on both these categories (OpenAI, 2024b). It is not clear if future AI systems will maintain this trend, but in light of recent progress, it may be prudent for nation-states to build up capacity both to track the national security threats that AI systems pose and to execute countermeasures to neutralize these threats. Against this backdrop of growing AI capabilities, U.S. federal legislators have started proposing nascent variations of ‘AI incident reporting regimes’ (118th Congress, 2024; Ross, 2024).
@@
We combine these two strands – increased government interest in both the national security risks posed by AI systems, and in AI incident regimes – to focus on the role that a legally mandated post-public deployment AI incident regime could play in giving a government agency oversight into national security threats posed by AI systems. Our ambition is to enable a government agency to maintain a comprehensive awareness of AI threats and rapidly counter any resulting risks to national security.
@@
In this paper, we put forward a proposal for an AI incident regime that will help to counter threats to national security posed by AI systems. We start the paper by introducing a novel term – ‘security-critical’ – to describe sectors that pose extreme risks to national security, before arguing that civilian nuclear power, aviation, and life sciences dual-use research of concern, and frontier AI development should all be considered ‘security-critical’ sectors. We then spend much of the paper detailing our proposal for an AI incident regime that aims to counter national security threats from AI systems. We justify each component of our AI incident proposal by demonstrating that it mirrors existing regimes in security-critical industries, specifically U.S. domestic incident regimes in civilian aviation, life sciences dual-use research of concern (e.g., genetic engineering), and nuclear power.
@@
Informed by this comparison to other high-risk areas, our proposed AI incident regime consists of three phases. First, a ‘Preparatory Phase, which details steps that AI providers should take in order to define which events count as ‘incidents’ ahead of time; then the ‘Rapid Response Phase,’ in which AI providers promptly notify a government agency about incidents, and this government agency taking steps to contain the incidents; and finally the ‘Hardening Defenses Phase,’ wherein AI providers improve their security and safety procedures in order to counter future threats to national security. We end our paper with a sketch of how our proposal would deal with a hypothetical spear-phishing attack.
@@
In summary, the core contribution of this paper is a proposal for a national security-focused AI incident regime [^1] We note that all references to an AI incident regime in this paper are tied to an AI incident regime specifically for national security threats, rather than other harms. which has three phases, with each component touching on either AI providers or the government agency :
@@
      AI providers should create ‘national security cases’ before publicly deploying frontier AI systems.
@@
      ‘Incident’ should be operationalized as ‘any event which weakens a claim made in a national security case for the AI system.’
@@
      Providers of frontier AI systems must notify a government agency immediately and no later than 24 hours after discovering an incident has occurred [^2] We define frontier AI systems as ’highly capable general-purpose AI models or systems that can perform a wide variety of tasks and match or exceed the capabilities present in the most advanced models’ (UK Government, 2024). We believe targeting frontier models ensures the most severe threats to national security can be rapidly identified and neutralized while still enabling broad economic benefits from the widespread adoption of less advanced AI systems..
@@
      If appropriate, the government agency shall coordinate an emergency containment to limit the harm caused by the incident.
@@
      The government agency has the ‘intelligence-gathering’ authority to access information that allows them to investigate the cause of an incident (consisting, for example, of extensive black-box access, some white-box access, and relevant documentation and technical results [^3] Note that it is not currently possible to use interpretability tools to reliably obtain a ‘decision rationale’ that ‘explains’ an AI systems output. Instead, we are recommending that a government agency has some white-box access to the AI system they are investigating in order to speed up their black-box investigation.).
@@
Our proposal for an AI incident regime is limited in scope: we detail an AI incident regime focussed purely on countering the most extreme threats to national security (see Section 2 for more justification) given the threat to people’s lives and livelihoods (Gaibulloev & Sandler, 2019; Iftikhar, 2024). Further, our proposal is designed to be agile, imposing a minimal burden on providers of AI systems that do not pose threats to national security. However, if AI systems start to possess dangerous capabilities, the regime ratchets upwards, offering ample intelligence-gathering and security-strengthening capabilities to a government agency and thus helping to counter potential national security threats from AI systems.
@@
2 Frontier AI development, nuclear power, aviation, and life sciences dual-use research of concern are all ‘security-critical’ sectors.
@@
In this Section, we coin the term ‘security-critical’ to describe sectors that pose extreme risks to national security. We argue that AI and all three of our selected case study sectors – nuclear power, aviation, and life sciences DURC – are ‘security-critical’ sectors. This leads us to the claim that incident regimes in these security-critical industries are likely good templates for the design of an AI incident regime: a claim we will use later on to justify each component of our AI incident regime proposal. However, we spend most of this Section demonstrating that there is a similarity in the threats to national security posed by AI systems and threats posed in nuclear power, aviation, and life sciences DURC in order to argue that it is natural to bucket all of these sectors together as ‘security-critical.’ We make this argument by first detailing the potential extreme threats posed in nuclear power, aviation, and life sciences DURC and then detailing the potential extreme threats posed by AI systems, showing that in the worst-case scenarios, there could be (at least) thousands of fatalities in a short space of time – making them all ‘security-critical.’
@@
We start by detailing the most extreme threats posed in nuclear power, aviation, and life sciences, demonstrating that these can lead to (at least) thousands of fatalities within months. For example, a 1975 NRC report on extreme risks posed by nuclear power offers an upper bound of the worst scenario, at 3000 fatalities “in a short time period after the incident” and up to 60,000 indirect fatalities from cancer (Rasmussen, 1975). Turning to historical precedents in nuclear power, the Chernobyl incident led to the deaths of 2 workers within hours of the plant meltdown and 22 within the following month due to the exposure to radiation (United Nations Scientific Committee, 2008), while one estimate from the International Atomic Energy Authority put the indirect number of fatalities due to cancer at 4000 (The Chernobyl Forum, 2006). Looking at historical incidents in aviation, the 1977 Tenerife collision between two aircraft resulted in the deaths of 583 people (Register-Guard, 1977), while the 9/11 attacks resulted in around 3,000 casualties (Encyclopedia Britannica, 2024). Meanwhile, in life sciences, one paper estimates that a potential pandemic pathogen escaping from a lab could lead to tens of millions of fatalities (Klotz & Sylvester, 2014). Historical precedent is demonstrated by events such as the 1979 anthrax lab leak in the Soviet Union, which led to 68 fatalities within 6 weeks of the leak occurring (Meselson et al., 1994). Meanwhile, COVID-19 is estimated to have resulted in over 7 million casualties as of January 2025 (World Health Organization, 2025), giving some sense of the possible worst-case consequences of a life sciences DURC lab leak.
@@
Now, we demonstrate that the most extreme national security threats from AI could also lead to at least thousands of fatalities within months of systems being deployed by detailing some of the extreme threats to national security that experts have cautioned about. We then explain that AI systems do not currently pose such threats, but note that if recent trends in AI capability growth continue, such threats will arise in the near future.
@@
First, we detail the worst-case national security threats that AI systems could pose. For example, some experts worry that AI systems could uplift the ability of malicious actors to create bioweapons, (Drexel & Withers, 2024; Reuters, 2023), which could cause a pandemic and lead to fatalities within weeks of initial infection (Baud et al., 2020). Alternatively, malicious actors could use AI systems to help with vulnerability discovery and exploitation for a large-scale cyber attack on critical national infrastructure (Lee et al., 2016; OECD, 2024). Such attacks could, for example, bring down the electricity grid within hours of being executed (Crowdstrike, 2024; Allianz Commercial, 2016). More speculatively, there is also the threat of loss of control of autonomous general-purpose AI systems (Bengio et al., 2025; Hendrycks et al., 2023; OECD, 2024), in which highly capable AI systems end up misaligned with the intentions of the AI provider (Bengio et al., 2025; Ngo et al., 2025). In this scenario, a misaligned, highly capable AI system could threaten national security, e.g., by creating bioweapons or executing a large-scale cyber attack (Bengio et al., 2024; Hendrycks et al., 2023).
@@
Next, we note a worrying trend in AI capabilities development that points towards AI systems posing threats to national security through cyber, bio, or loss of control threats within the next few years. Despite the fact that as of March 2025, publicly deployed AI systems do not appear to pose much danger (Bengio et al., 2025; Christopher A. Mouton, 2023), recent growth in general AI capabilities has been fast (Maslej et al., 2024; Roser, 2022), and more recently there has been growth in capabilities that pose threats to national security. Regarding the former, Anthropic CEO Dario Amodei has said that within 3 years, we will have a “country of geniuses in a datacenter” (Amodei, 2024) and that “AI could surpass almost all humans at almost everything” (Edwards, 2025). Further, a recent report co-authored by 96 world-leading AI experts includes a Chair’s Note from Yoshua Bengio claiming that recent evidence points towards “the pace of advances in AI capabilities … remain[ing] high or even accelerat[ing]”(Bengio et al., 2025). Against this backdrop of general AI capabilities increasing, there has also been fast progress in dangerous capabilities that pose threats to national security: a leading AI system from August 2024 (GPT-4o) was rated as posing a “Low” risk on all of these domains (OpenAI, 2024a) while just 7 months later a frontier system was rated as posing a Medium risk on all these domains (OpenAI’s Deep Research, released in January 2025) (OpenAI, 2025). If this trend is maintained, AI systems will soon pose threats to national security on par with those from nuclear power, aviation, and life sciences DURC.
@@
Hence, in this sub-section, we demonstrated that AI systems may soon pose national security threats similar to those in nuclear power, life sciences DURC, and aviation, involving thousands of fatalities in a short space of time. Given the similarity in threats, we claim that incident regimes in these security-critical sectors are promising templates for an AI incident regime – and this is a claim that will underpin each part of our AI incident regime proposal.
3 Our proposal for an AI incident regime
@@
Our novel proposal for an AI incident regime consists of three phases, with each being backed up by our review of incident regimes in security-critical sectors: a ‘Preparatory Phase’ where AI providers pre-determine ahead of time what events count as incidents; a ‘Rapid Response Phase’ wherein AI providers promptly notify a government agency about incidents, and the government agency then contains the incident; and a ‘Hardening Defenses Phase,’ in which the government agency takes steps to counter future national security threats.
@@
We take most of this Section (3.1-3.3) to articulate in detail each of these phases of our AI incident regime proposal. We justify, in turn, each component of our proposed AI incident regime by demonstrating that the component mirrors incident regimes in other security-critical sectors. In particular, we make the comparison between our AI incident regime proposal and: in nuclear power, the incident regime run by the Nuclear Regulatory Commission (NRC); in life sciences, the National Institute for Health’s (NIH) incident regime on genetic engineering and the Federal Select Agents Program (FSAP) incident regime on dangerous biological agents; and in aviation, the National Transportation Safety Board’s (NTSB) incident investigations and the Federal Aviation Authority’s (FAA) post-incident rule-making. We then end the paper with a hypothetical scenario tying our blueprint together in Section 4. Before detailing each phase of our AI incident regime, we present a short summary, detailing for each phase the goal and how we expect AI providers or the government agency to act [^4] The components in our incident regime proposal should not be considered sufficient for an AI incident regime to adequately counter potential threats to national security. For example, in order for the incident regime to be successful, it may be important to set up infrastructure to allow consumers or downstream developers to report incidents (in addition to AI providers) in order to ensure that the regime comprehensively catches incidents.:
@@
      AI providers should report as an incident any event that weakens a claim made in a ‘national security case’ for the AI system
@@
      Providers of frontier AI systems must notify a government agency immediately, and no later than 24 hours, after they discover that an incident has occurred [^5] We define frontier AI systems as “highly capable general-purpose AI models or systems that can perform a wide variety of tasks and match or exceed the capabilities present in the most advanced models” (UK Government, 2024). We believe targeting frontier models ensures the most severe threats to national security can be rapidly identified and neutralized while still enabling broad economic benefits from the widespread adoption of less advanced AI systems..
@@
      If appropriate, the government agency will coordinate an emergency containment to limit the harm caused by the incident
@@
      The government agency has the ‘intelligence-gathering’ authority to perform on incidents through extensive access to the potentially compromised AI system (consisting, for example, of extensive black-box access, some white-box access, and relevant documentation and technical results [^6] Note that it is not currently possible to use interpretability tools to reliably obtain a ‘decision rationale’ that ‘explains’ an AI systems output. Instead, we are recommending that a government agency has some white-box access to the AI system they are investigating in order to speed up their black-box investigation).
@@
      A government agency has a ‘security-strengthening authority’ to make recommendations as to how AI providers’ safety and security infrastructure should be modified based on the results of their root cause analyses. Under this authority, all relevant AI providers must implement these recommendations in order to counter future threats to national security.
@@
3.1 ‘Preparatory Phase’: AI providers should create national security safety cases pre-deployment and should report as an incident ‘any event which weakens a claim made in the national security case.’
@@
The first phase of our proposal for an AI incident regime – featuring a novel operationalization of ‘AI incident’ – consists of two preparatory steps that AI providers must take in order to quickly report an incident when it arises: (i) AI providers must create national security cases for their frontier AI systems before deploying them publicly and (ii) AI providers must report as an incident ‘any event which weakens the evidence for a claim or argument chain made in an AI system’s national security case.’ We start this Section by detailing how an AI provider could comply with each of these with each of these steps, and then go on to justify our proposal for a ‘Preparatory Phase’ with a three-part argument: first, we show that incident regimes in other security-critical sectors have both clear criteria for what counts as an ‘incident,’ and include near-misses or precursory harms as incidents. Then, we detail how an existing operationalization of AI incidents in the literature fails to fulfill these criteria before finally showing that our proposal addresses these shortcomings.
@@
We start by presenting the first obligation for AI providers under the ‘Preparatory Phase’: AI providers should produce a ‘national security safety case’ before publicly deploying any frontier models [^7] This paper focuses on AI incidents caused by AI systems that have been publicly deployed. In principle, national security cases could be written before public deployment in order to track national security incidents posed by AI providers’ internal deployment of frontier AI systems.. A ‘national security case’ is a security and safety framework adapted from previous frameworks in the literature. It is a document laying out a structured argument making the case that a given AI system does not pose national security risks in public deployment (Ministry of Defence, 2017)– perhaps because it does not pose an unacceptable increase in biorisk, or risk of a large-scale cyber attack or risk of loss of control of an autonomous AI system [^8] The ‘national security case’ is an adaptation of a previous AI governance framework – AI safety cases – for a national security setting (Balesni et al., 2024; Clymer et al., 2024; Buhl et al., 2024; Goemans et al., 2024; METR, 2024; UK Government, 2024). This structured argument comprises a series of claims, each of which needs to be supported by verifiable evidence about the AI system in question. [^9] It is outside the scope of this paper to make recommendations on which threats an AI national security case should focus on, and therefore this is a promising question for further research. In addition to the cyber, bio, and loss of control threats already mentioned, some other plausible threats include mass persuasion and threats to the financial system from automated trading.. For example, an AI provider might make the high-level claim in their ‘national security case’ that a given AI system does not pose an unacceptable cyber threat. The AI provider might justify this high-level claim, as in Figure 1, by a series of more concrete claims that the AI systems cannot help certain actors perform certain kinds of cyber attack – for example, claiming that the AI system cannot assist a technical user who is not a cybersecurity expert in discovering and exploiting cyber vulnerabilities in a realistic setting (C3.1). The AI provider then justifies these sub-claims with a body of empirical evidence about the AI system, usually dangerous capability or propensity, evaluation test results – for example, demonstrating through a ‘human uplift’ study that, when paired with AI systems, a human novice does not improve at cyber vulnerability discovery, compared to doing it by themselves. For more concrete details on what an AI ‘national security case’ might look like in the cyber domain, see (Goemans et al., 2024).
@@
Figure 1: A visualization of part of a ‘national security case’ which argues that a given AI system does not pose a cyber threat (taken from (Goemans et al., 2024)).